OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13 Answer x + 5y + 6 –x – 5y + 6 x + 5y – 6 –x – 5y – 6

OpenStudy (anonymous):

if ur gonna substitute the z into the 2nd and third equation u want to solve for z in the first equation and whatever z equals is what u substitue: x+5y-z=6 x+5y-z+z=6+z x+5y=6+z x+5y-6=z+6-6 z=x+5y-6 so u would substitute x+5y-6 into the other equations

OpenStudy (anonymous):

for z

OpenStudy (anonymous):

x+5y-6 is ur answer to be clear btw